what is the first step to this equation
8x-9y=19
4x+y=-7

Found 2 solutions by TutorDelphia, stanbon:
Answer by TutorDelphia(193)

The first step to solving it? I'd multiple the second equation by -2 (and then add the two equations together causing the x term to drop out, solve for y and then plug in the value I got for y into one of the two initial equations)

Answer by stanbon(75887)

what is the first step to this equation
8x-9y=19
4x+y=-7
-----
Multiply thru the bottom equation by 2 to get:
8x-9y=19
8x+2y=-14
==============
Cheers,
Stan H.
================